# RealSR-R1: Reinforcement Learning for Real-World Image Super-Resolution with Vision-Language Chain-of-Thought

### v2

@@ Abstract
Real-World Image Super-Resolution is one of the most challenging task in image restoration. However, existing methods struggle with an accurate understanding of degraded image content, leading to reconstructed results that are both low-fidelity and unnatural. We present RealSR-R1 in this work, which empowers the RealSR models with understanding and reasoning capabilities. Inspired by the success of Chain of Thought (CoT) in large language models (LLMs), we simulate the human process of handling degraded images and propose the VLCoT framework, which integrates vision and language reasoning. The framework aims to precisely restore image details by progressively generating more comprehensive text and higher-resolution images. To overcome the challenge of traditional supervised learning CoT failing to generalize to real-world scenarios, we introduce, for the first time, Group Relative Policy Optimization (GRPO) into the Real-World Image Super-Resolution task. We propose VLCoT-GRPO as a solution, which designs four reward functions: (1) Format reward, used to standardize the CoT process; (2) Degradation reward, to incentivize accurate degradation estimation; (3) Understanding reward, to ensure the accuracy of the generated content; and (4) Generation reward, where we propose using a visual expert model to evaluate the quality of generated images, encouraging the model to generate more realistic images. Extensive experiments demonstrate that our proposed RealSR-R1 can generate realistic details and accurately understand image content, particularly in semantically rich scenes or images with severe degradation. Code: https://github.com/Junboooo/RealSR-R1.
@@ 1 Introduction
Figure 1: (a)The GAN and diffusion methods themselves do not possess the capability to understand the content of images. (b) Directly interpreting and generating detailed descriptions of degraded images leads to incorrect understanding and image restoration. (c)RealSR-R1 simulates the human image restoration process by progressively refining the understanding of image content and generating higher-quality images.
@@
Recently, Multimodal Large Language Model (MLLM) leverages high-capacity networks and large-scale data training, demonstrating preliminary image understanding and generation capabilities in RealSR tasks [15]. Despite the stronger understanding capabilities of such methods, directly providing detailed descriptions of severely degraded and semantically ambiguous scenarios leads to content misinterpretation and detail confusion, resulting in artifacts, random generation, or structurally inconsistent outputs. As illustrated in Fig. 1(b), the severely degraded LR depicts a scene of rooftops and sky, yet the MLLM misinterprets this scene as "the building appears to be a church, as it has a steeple and a cross on top," ultimately leading to the generation of hallucinated images.
In natural language processing (NLP) and MLLM, researchers have observed that directly mapping inputs to outputs on complex tasks often leads to biased comprehension or suboptimal outcomes [16]. Consequently, the Chain-of-Thought (CoT) mechanism [16] is proposed, guiding models to simulate the human thinking process by explicitly constructing intermediate reasoning steps. CoT significantly enhances model understanding and output reliability [17, 18, 19, 20]. Inspired by this mechanism, we attempt to introduce CoT into RealSR to address the complex restoration problem. However, existing CoT approaches in vision typically involve intermediate reasoning processes in either pure language [21, 22, 23] or pure image [24] form. In contrast, for RealSR tasks, CoT not only requires reasoning expressed in textual form for understanding, but also needs reasoning at the visual level to accurately capture structural and perceptual details. To design a vision-language CoT process suitable for the RealSR task, we reflect on the human restoration process. We observe that it is a process where generation and understanding mutually reinforce each other, relying on extensive world priors. Specifically, humans first assess the degree of image degradation and roughly understand the image content (rather than directly jumping to pixel-level reconstruction). Then, the image is progressively restored in a coarse-to-fine manner, with a continuous deepening of understanding about the image content throughout the process.
Motivated by these insights, we propose VLCoT, a multi-step reasoning process that combines vision and language in a coarse-to-fine manner. As shown in Fig. 1(c), for the input LR, the model first estimates the degradation degree, coarse understanding, and coarse restored image, and then progressively generates more detailed textual understanding descriptions and higher-resolution images. Ultimately, the proposed VLCoT accurately understands the brick walls and roofs in the image and achieves the most precise image restoration.
@@
Furthermore, we propose VLCoT-GRPO, which leverages Group Relative Policy Optimization (GRPO) [25] to enhance and optimize the image-text integrated reasoning process of VLCoT. We introduce reinforcement learning (RL) for two reasons: (1) Recent MLLMs have already demonstrated basic CoT reasoning capabilities in individual modalities such as text or image. To address this limitation, we aim to foster active exploration that unlocks synergistic reasoning across modalities. (2) Supervised fine-tuning (SFT) relies on imitating high-quality reasoning paths, which may lead the model to adopt "seemingly correct" patterns while failing to identify and differentiate flawed reasoning effectively. In contrast, RL encourages exploration of diverse reasoning paths, strengthens correct reasoning patterns through reward signals, and simultaneously identifies and suppresses incorrect reasoning patterns. This approach is better suited for non-regular problems like RealSR.
Unlike standard NLP tasks that typically employ clear-cut reward signals, RealSR lacks a unified, standardized reward function, posing challenges for RL-based optimization. To tackle these issues, we propose four reward functions to comprehensively guide the VLCoT generation process: (1) Format Reward: Enforces structural conventions to ensure a well-formatted, hierarchical generation process. (2) Degradation Reward: Ensure accurate perception of degradation. (3) Understand Reward: Assesses the model’s accuracy in understanding the LR image contents, thus ensuring reasonableness and precision of the outputs. (4) Generation Reward: Using a powerful prior visual expert model, a multi-faceted evaluation of the image quality, realism and consistency of a group of outputs is provided.
Building on the aforementioned reasoning and training method, we propose RealSR-R1, the first reasoning-enhanced RealSR model. Experimental results demonstrate the effectiveness of our approach, and qualitative analyses further indicate that RealSR-R1 produces images more aligned with human preferences and demonstrates higher robustness when handling complex degradations or intricate scenes.
@@
  • We propose RealSR-R1, the first reasoning-enhanced solution to the RealSR problem for more stable and trustworthy image reconstruction.
  • We propose VLCoT that establishes a progressive, multi-step reasoning process and combines vision and language, mimicking the process of humans’ transition from understanding to image reconstruction, gradually refining the image from coarse to fine.
  • We present VLCoT-GRPO, a reinforcement learning method for RealSR, which effectively optimizes the reasoning process of VLCoT by leveraging multiple carefully designed rewards.
@@ 2 Related Work
The o1 [34] model proposed by OpenAI has attracted widespread attention, particularly for its advancements in enhancing reasoning capabilities through reinforcement learning (RL). DeepSeek-R1 [35] introduced a rule-based reward mechanism, Group Relative Policy Optimization (GRPO), a training paradigm that requires comprehensive reasoning before final answer generation. In recent years, this methodology has been widely adopted in the development of MLLMs [36, 37, 38, 39, 40] with task-specific reward functions, such as correctness [22] or IoU [41]. However, most of these tasks primarily apply reinforcement learning to pure text [21, 22, 23] or image [24, 42] inference processes. There remains a significant research gap in the RealSR tasks, which require both generation and comprehension. To address this gap, our work is the first to explore reinforcement learning for large reasoning models in the image restoration field.
@@ 3.1 Preliminary
$\mathcal{J}_{GRPO}(\theta)=\mathbb{E}\left[(q,a)\sim D,\left\{o_{i}\right\}_{i=1}^{G}\sim\pi_{\theta_{\text{old }}}(O\mid q)\right]$ (1)
$\frac{1}{G}\sum_{i=1}^{G}\left(\min\left(\frac{\pi_{\theta}\left(o_{i}\mid q\right)}{\pi_{\theta_{\text{old }}}\left(o_{i}\mid q\right)}A_{i},\operatorname{clip}\left(\frac{\pi_{\theta}\left(o_{i}\mid q\right)}{\pi_{\theta_{\text{old }}}\left(o_{i}\mid q\right)},1-\varepsilon,1+\varepsilon\right)A_{i}\right)-\beta\mathbb{D}_{KL}\left(\pi_{\theta}||\pi_{ref}\right)\right).$
@@ 3.2 VLCoT
VLCoT process design.
Unlike traditional single-modal Chain-of-Thought design, human image restoration does not separate the processes of understanding and generation. Therefore, we propose Vision-Language Chain of Thought (VLCoT), a coarse-to-fine, multi-step reasoning process that integrates both textual and visual information. As illustrated in Fig. 2, given an LR image, we guide the model to generate VLCoT reasoning with the following prompt: ‘Perceive the degradation, understand the image content, and restore the high-quality image step by step (simulating the image restoration process in three steps from coarse to fine).’ Then VLCoT decomposes the image restoration process into a sequence of multimodal reasoning steps, including: (1) degradation perception step; (2) coarse restoration step (coarse understanding and image generation); (3) middle-detail restoration step (middle-detail understanding and image generation); and (4) fine-detail restoration step (fine-detail understanding and image generation). This design mimics the human perception-to-reasoning pathway, starting from degradation awareness, followed by a coarse-to-fine process of understanding and image generation, enabling the model to progressively and robustly restore both semantic and structural information from severely degraded images.
Cold-start Initialization for VLCoT.
To ensure that the model has stable output style and basic image restoration capabilities from the start, we perform cold-start training on VLCoT before fine-tuning the entire model using reinforcement learning. Specifically, we first construct paired input-output and CoT datasets, which will be detailed in the following sections, and process the concatenated text-image token sequences through autoregressive learning. Given a multimodal sequence $X=\left(x_{1},x_{2},\ldots,x_{T}\right)$, where each $x_{t}$ may be a text token, image patch, or other modality embedding, the learning objective is to maximize the likelihood of the next element conditioned on previous observations: $P_{\left(x\mid y\right)}(\theta)=-\sum_{t=1}^{T}\log P_{\theta}\left(x_{t}\mid x_{<t},I\right)$ where I is the input and $x_{<t}=\left(x_{1},\ldots,x_{t-1}\right)$ represents the preceding context. The overall training optimization procedure minimizes the stepwise Cross-Entropy loss between the predictions and the targets:
@@
Degradation perception step.
In the degradation perception step, we follow [6, 10] by randomly selecting degradation (e.g., Gaussian or Poisson noise) and sampling all parameters (e.g., noise intensity $\mu_{1}\in[a,b]$) from a uniform distribution to determine the degradation. However, these degradation ground truth (GT) values with specific parameters are difficult for MLLMs to comprehend. Therefore, we further transform them into user-friendly summary text representations. Specifically, we follow [45, 46] to discretize the sampling distribution of the parameters for each degradation component (e.g., noise, blur, JPEG artifacts) by evenly dividing it into discrete intervals. These intervals are then discretely described and summarized to represent the degradation level. For example, we divide the distribution of noise level $\varphi_{1}$ into three uniform intervals, which are described as "light," "medium," and "heavy." Finally, the overall degradation representation combines the descriptions of all components, such as "[deblurring description, noise description,…, resizing description]".
Restoration step.
In the image generation stage of each restoration step, the ground-truth supervision for the coarse, middle-detail, and fine-detail images is derived from the HQ image downsampled by factors of 4 and 2, as well as the original resolution, respectively. For the understanding text generation, we utilize Qwen2.5-VL [47] to produce descriptions of the GT images at each level (coarse, middle-detail, and fine-detail restoration images) through a three-round dialogue process. The detailed prompts used in the dialogue rounds are provided in the Appendix A.2.
Tokenization.
For RealSR tasks, high compression ratios and limited codebook capacity result in significant information loss during the extraction of discrete image representations. To mitigate this issue, we adopt the VQGAN tokenizer from LlamaGen [48], which offers a lower downsampling factor compared to the original Lumina-mGPT (e.g., 8×8$\times$8 × vs. 16×16$\times$16 ×) and a larger visual vocabulary size (e.g., 16,536 vs. 8,192). This modification substantially reduces information loss during the encoding process, thereby improving the quality of reconstructed images in the Real-ISR task. For text tokenization, we retain the BPE tokenizer used in Lumina-mGPT, which originally has a vocabulary size of 65,536. To accommodate the expanded image tokenizer, the BPE tokenizer’s vocabulary size is increased to 73,728, enabling it to handle a larger set of discrete image tokens.

### v3

@@ Abstract
Real-World Image Super-Resolution is one of the most challenging task in image restoration. However, existing methods struggle with an accurate understanding of degraded image content, leading to reconstructed results that are both low-fidelity and unnatural. We present RealSR-R1 in this work, which empowers the RealSR models with understanding and reasoning capabilities. Inspired by the success of Chain of Thought (CoT) in large language models (LLMs), we simulate the human process of handling degraded images and propose the VLCoT framework, which integrates vision and language reasoning. The framework aims to precisely restore image details by progressively generating more comprehensive text and higher-resolution images. To overcome the challenge of traditional supervised learning CoT failing to generalize to real-world scenarios, we introduce, for the first time, Group Relative Policy Optimization (GRPO) into the Real-World Image Super-Resolution task. We propose VLCoT-GRPO as a solution, which designs four reward functions: (1) Format reward, used to standardize the CoT process; (2) Degradation reward, to incentivize accurate degradation estimation; (3) Understanding reward, to ensure the accuracy of the generated content; and (4) Generation reward, where we propose using a visual expert model to evaluate the quality of generated images, encouraging the model to generate more realistic images. Extensive experiments demonstrate that our proposed RealSR-R1 can generate realistic details and accurately understand image content, particularly in semantically rich scenes or images with severe degradation. Our source codes have been released at https://github.com/Junboooo/RealSR-R1
Index Terms:
Real-World Image Super-Resolution, Reinforcement Learning, Chain of Thought, Multimodal Large Language Model.
Figure 1: (a) The GAN and diffusion methods themselves do not possess the capability to understand the content of images. (b) Directly interpreting and generating detailed descriptions of degraded images leads to incorrect understanding and image restoration. (c) RealSR-R1 simulates the human image restoration process by progressively refining the understanding of image content and generating higher-quality images. (d) Incorrect semantics lead the model to reconstruct erroneous content, whereas correct semantic recovery of fine-grained details, such as plant textures, is crucial.
@@ 1 Introduction
Recently, Multimodal Large Language Model (MLLM) leverages high-capacity networks and large-scale data training, demonstrating preliminary image understanding and generation capabilities in RealSR tasks [59]. However, we observe that in Fig. 1(b), the severely degraded LR depicts a scene of rooftops and sky, yet the MLLM misinterprets this scene as ”the building appears to be a church, as it has a steeple and a cross on top,” consequently producing hallucinated structures in the reconstructed image. This phenomenon exposes a fundamental issue in directly applying the semantics-driven generation logic of MLLMs to image reconstruction tasks: when confronted with heavily degraded LR inputs, the model may fail to correctly interpret the visual content and instead rely on its internal, generalized semantic priors to “hallucinate” plausible but incorrect details.
To further validate the impact of semantic understanding correctness on reconstruction quality, we design a controlled experiment based on the PURE [59] by explicitly manipulating the correctness of semantic understanding (incorrect, none, and correct). As shown in Fig. 1 (d), incorrect semantic understanding leads the model to generate erroneous structural content, whereas only accurate semantic interpretation enables the recovery of fine-grained details such as plant textures. These observations raise a key question: how can a model acquire accurate semantic understanding when the input is severely degraded?
For this question, we first reflect on how humans perform visual restoration. We observe that semantic understanding and image reconstruction are not independent processes but are inherently interdependent [52, 33, 21]. Specifically, humans typically begin by assessing the degree of degradation and forming a coarse semantic understanding of the scene, rather than attempting pixel-level restoration directly. The reconstruction process then proceeds in a coarse-to-fine manner, during which semantic understanding is progressively refined as more visual details are recovered. This iterative interaction ultimately yields both a visually detailed reconstruction and an accurate semantic interpretation. Therefore, establishing a bidirectional reasoning loop between visual restoration and semantic understanding is crucial.
Motivated by these insights, we propose the vision-language Chain-of-Thought (VLCoT), a multi-step vision-language reasoning framework that models the mutual reinforcement between semantic understanding and image reconstruction. Instead of performing one-shot restoration based on uncertain semantics, VLCoT adopts a coarse-to-fine reasoning paradigm, where vision and language iteratively refine each other. As shown in Fig. 1(c), given a degraded LR image, the model first produces an initial coarse understanding along with a coarse reconstructed image. These intermediate results then serve as improved visual and semantic cues for subsequent reasoning steps, progressively enhancing both the textual understanding and image fidelity. Ultimately, VLCoT achieves accurate recognition of fine-grained structures—such as the brick walls and roof—and produces the most faithful high-resolution restoration.
@@
While VLCoT establishes a mutual interaction between semantic understanding and visual reconstruction, its reasoning quality still depends heavily on supervised learning, which limits the model’s ability to self-correct and generalize. To further enhance this process, we propose VLCoT-GRPO, which adapts Group Relative Policy Optimization (GRPO) [48] to optimize the image-text integrated reasoning of VLCoT through reinforcement learning (RL). We introduce RL for two main reasons. (1) Although recent MLLMs exhibit basic chain-of-thought (CoT) reasoning within individual modalities (e.g., text or image), they struggle to perform synergistic reasoning across modalities. To address this limitation, we aim to foster active exploration that unlocks synergistic reasoning across modalities. (2) Supervised fine-tuning (SFT) relies on imitating high-quality reasoning paths, and this approach risks overfitting to “seemingly correct” but suboptimal patterns; in contrast, RL provides reward-driven feedback that strengthens genuinely correct reasoning behaviors while penalizing erroneous ones. This approach is better suited for non-regular problems like RealSR.
Unlike standard NLP or reasoning tasks where explicit reward signals (e.g., accuracy or token-level correctness) are readily available, RealSR lacks a unified and standardized reward metric, making RL-based optimization particularly challenging. The quality of restoration involves both semantic fidelity and visual realism, which cannot be captured by a single scalar reward. To address this issue, we design four complementary reward functions to comprehensively guide the vision-language reasoning process of VLCoT-GRPO: (1) Format Reward: enforces structural and hierarchical generation conventions, ensuring a consistent reasoning format; (2) Degradation Reward: evaluates whether the model correctly perceives the degradation level, which is critical for guiding restoration difficulty; (3) Understanding Reward: measures the semantic accuracy of the model’s interpretation of LR content, ensuring logical and perceptual coherence; and (4) Generation Reward: employs a powerful pretrained visual expert to provide multi-dimensional assessments of image realism, quality, and consistency across outputs. These complementary rewards establish a holistic feedback mechanism that drives VLCoT-GRPO to achieve more stable training, stronger semantic-visual alignment, and superior reconstruction performance under severe degradations.
Building on the proposed vision-language reasoning framework and its reinforcement optimization strategy, we propose RealSR-R1, the first reasoning-enhanced RealSR model. Benefiting from the multi-level reward guidance and cross-modal reasoning refinement, RealSR-R1 not only achieves superior quantitative performance but also produces perceptually realistic results that align more closely with human visual preferences. Extensive experiments further demonstrate its robustness in handling challenging degradations and semantically intricate scenes, marking a significant step toward reasoning-aware RealSR.
@@
  • We propose RealSR-R1, the first reasoning-enhanced solution to the RealSR problem, highlighting the importance of correct semantic understanding for reconstruction to achieve more stable and trustworthy image reconstruction.
  • We propose VLCoT, a progressive multi-step vision-language reasoning framework that emulates the human process of “understanding before reconstructing.” VLCoT jointly refines semantic understanding and visual restoration in a coarse-to-fine manner.
  • We present VLCoT-GRPO, a reinforcement learning framework for multi-modal joint reasoning, which aligns and optimizes visual and textual reasoning with multiple reward signals, enhancing semantic-visual coherence and reasoning robustness.
@@ 2 Related Work
2.2 Chain-of-Thought.
In natural language processing (NLP) and MLLM, researchers have observed that directly mapping inputs to outputs on complex tasks often leads to biased comprehension or suboptimal outcomes [60]. Consequently, the Chain-of-Thought (CoT) mechanism [60, 25, 22, 75] is proposed, guiding models to simulate the human thinking process by explicitly constructing intermediate reasoning steps. CoT significantly enhances model understanding and output reliability [68, 38, 39, 80]. However, existing CoT approaches in vision and vision-language domains typically perform reasoning within a single modality, relying solely on textual descriptions [47, 35, 40] or visual features [16]. Such designs are insufficient for Real Image Super-Resolution (RealSR), where both semantic understanding and visual reconstruction are inherently intertwined. Motivated by the mutual reinforcement between understanding and reconstruction, we propose the Vision-Language Chain-of-Thought (VLCoT) framework tailored for RealSR. VLCoT adopts a coarse-to-fine, alternating reasoning paradigm, where semantic reasoning provides guidance for visual restoration, and intermediate reconstruction results in turn refine semantic understanding. This bidirectional reasoning process enables the model to progressively align semantics and perception, achieving coherent and faithful image reconstruction even under severe degradations.
@@
Reinforcement learning (RL) has recently gained significant attention for its effectiveness in enhancing reasoning capabilities of large models. The o1 [20] model proposed by OpenAI has attracted widespread attention, particularly for its advancements in enhancing reasoning capabilities through RL. DeepSeek-R1 [15] introduced a rule-based reward mechanism, Group Relative Policy Optimization (GRPO), a training paradigm that requires comprehensive reasoning before final answer generation. In recent years, this methodology has been widely adopted in the development of MLLMs [7, 37, 67, 72, 11] with task-specific reward functions, such as correctness [35] or IoU [34]. However, most of these tasks primarily apply reinforcement learning to pure text [47, 35, 40] or image [16, 4] inference processes. In contrast, RealSR requires joint optimization of semantic reasoning and visual reconstruction, which are tightly coupled yet rarely addressed under the RL framework. To bridge this gap, we explore RL for large reasoning models in image restoration and propose VLCoT-GRPO, a multi-modal RL framework that coordinates and optimizes both visual and textual reasoning for more coherent and faithful RealSR reconstruction.
@@ 3.1 Preliminary
$\mathcal{J}_{GRPO}(\theta)=\mathbb{E}\left[(q,a)\sim D,\ \{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(O\mid q)\right]$ (1)
$\frac{1}{G}\sum_{i=1}^{G}\Bigg(\min\Bigg(\frac{\pi_{\theta}(o_{i}\mid q)}{\pi_{\theta_{\text{old}}}(o_{i}\mid q)}A_{i},\operatorname{clip}\!\left(\frac{\pi_{\theta}(o_{i}\mid q)}{\pi_{\theta_{\text{old}}}(o_{i}\mid q)},\ 1-\varepsilon,\ 1+\varepsilon\right)A_{i}\Bigg)$
$\qquad-\beta\,\mathbb{D}_{KL}\!\left(\pi_{\theta}\,\|\,\pi_{\text{ref}}\right)\Bigg).$
@@ 3.2 VLCoT
VLCoT process design. Unlike traditional single-modal Chain-of-Thought design, human image restoration does not separate the processes of understanding and generation. Therefore, we propose Vision-Language Chain of Thought (VLCoT), a coarse-to-fine, multi-step reasoning process that integrates both textual and visual information. As illustrated in Fig. 2, given an LR image, we guide the model to generate VLCoT reasoning with the following prompt: ‘Perceive the degradation, understand the image content, and restore the high-quality image step by step (simulating the image restoration process in three steps from coarse to fine).’ Then VLCoT decomposes the image restoration process into a sequence of multimodal reasoning steps, including: (1) degradation perception step; (2) coarse restoration step (coarse understanding and image generation); (3) middle-detail restoration step (middle-detail understanding and image generation); and (4) fine-detail restoration step (fine-detail understanding and image generation). This design mimics the human perception-to-reasoning pathway, starting from degradation awareness, followed by a coarse-to-fine process of understanding and image generation, enabling the model to progressively and robustly restore both semantic and structural information from severely degraded images.
Cold-start Initialization for VLCoT. To ensure that the model has stable output style and basic image restoration capabilities from the start, we perform cold-start training on VLCoT before fine-tuning the entire model using reinforcement learning. Specifically, we first construct paired input-output and CoT datasets, which will be detailed in the following sections, and process the concatenated text-image token sequences through autoregressive learning. Given a multimodal sequence $X=\left(x_{1},x_{2},\ldots,x_{T}\right)$, where each $x_{t}$ may be a text token, image patch, or other modality embedding, the learning objective is to maximize the likelihood of the next element conditioned on previous observations: $P_{\left(x\mid y\right)}(\theta)=-\sum_{t=1}^{T}\log P_{\theta}\left(x_{t}\mid x_{<t},I\right)$ where I is the input and $x_{<t}=\left(x_{1},\ldots,x_{t-1}\right)$ represents the preceding context. The overall training optimization procedure minimizes the stepwise Cross-Entropy loss between the predictions and the targets:
@@
Degradation perception step. In the degradation perception step, we follow [55, 63] by randomly selecting degradation (e.g., Gaussian or Poisson noise) and sampling all parameters (e.g., noise intensity $\mu_{1}\in[a,b]$) from a uniform distribution to determine the degradation. However, these degradation ground truth (GT) values with specific parameters are difficult for MLLMs to comprehend. Therefore, we further transform them into user-friendly summary text representations. Specifically, we follow [77, 9] to discretize the sampling distribution of the parameters for each degradation component (e.g., noise, blur, JPEG artifacts) by evenly dividing it into discrete intervals. These intervals are then discretely described and summarized to represent the degradation level. For example, we divide the distribution of noise level $\varphi_{1}$ into three uniform intervals, which are described as ”light,” ”medium,” and ”heavy.” Finally, the overall degradation representation combines the descriptions of all components, such as ”[deblurring description, noise description,…, resizing description]”.
Restoration step. In the image generation stage of each restoration step, the ground-truth supervision for the coarse, middle-detail, and fine-detail images is derived from the HQ image downsampled by factors of 4 and 2, as well as the original resolution, respectively. For the understanding text generation, we utilize Qwen2.5-VL [2] to produce descriptions of the GT images at each level (coarse, middle-detail, and fine-detail restoration images) through a three-round dialogue process. The detailed prompts used in the dialogue rounds are provided in the Fig 3. Each round is guided by a specific designed prompts to elicit increasingly fine-grained textual understanding of the visual content.
Tokenization. For RealSR tasks, high compression ratios and limited codebook capacity result in significant information loss during the extraction of discrete image representations. To mitigate this issue, we adopt the VQGAN tokenizer from LlamaGen [51], which offers a lower downsampling factor compared to the original Lumina-mGPT (e.g., $8\times$ vs. $16\times$) and a larger visual vocabulary size (e.g., 16,536 vs. 8,192). This modification substantially reduces information loss during the encoding process, thereby improving the quality of reconstructed images in the Real-ISR task. For text tokenization, we retain the BPE tokenizer used in Lumina-mGPT, which originally has a vocabulary size of 65,536. To accommodate the expanded image tokenizer, the BPE tokenizer’s vocabulary size is increased to 73,728, enabling it to handle a larger set of discrete image tokens.